# KANELÉ: Kolmogorov–Arnold Networks for Efficient LUT-based Evaluation

### v1

@@ 5.4. Comparison with Prior KAN-FPGA Literature
We benchmarked KANELÉ against the KAN FPGA implementation by Tran et al. (Tran_KAN_bad) on the Moons, Wine, and Dry Bean datasets. The results, detailed in Table 4, demonstrate that KANELÉ offers a dramatic improvement in hardware efficiency and performance while maintaining or exceeding the accuracy of the previous work.
@@ 5.5. Comparison with hls4ml MLPerf Tiny
To understand the performance of KANELÉ on more complex datasets, we compare the framework to hls4ml (Borras_hls4ml) on the ToyADMOS dataset, part of the MLPerf Tiny benchmark suite. Other LUT-based neural networks have not attempted this benchmark, possibly due to its high complexity and/or non-classification-based training scheme. The results in Table 5 indicate that KANELÉ achieves substantially better performance than prior approaches in terms of both resource efficiency, latency, and power. These improvements highlight the potential of KANELÉ for future studies involving more complex datasets as well as tasks beyond classification. Specifically, KANELÉ eliminates the need for BRAM, LUTRAM, and DSPs, while reducing LUT usage by 41.7% and FF usage by 71.4% relative to hls4ml. In terms of performance, KANELÉ delivers $330\times$ higher throughput, $643\times$ lower latency and a $9,840\times$ reduction in energy per inference. These results establish KANELÉ as a highly efficient alternative to existing FPGA neural implementations, with strong potential for scaling to more complex datasets and tasks beyond classification.
@@ 5.7. Extension to Real-time Control Systems
To demonstrate that the KANELÉ paradigm extends well beyond the traditional supervised learning tasks typically studied in the LUT-based neural network community, we apply our framework to a reinforcement learning benchmark. Specifically, we evaluate on the HalfCheetah environment, a classic continuous control task in reinforcement learning (RL), most commonly accessed through the MuJoCo physics simulator via OpenAI Gym (now Gymnasium) (Mujoco). The objective is to learn a policy that enables a simulated two-legged agent to run as fast and as stably as possible. This environment is widely used as a standard benchmark to compare RL algorithms and function approximators. While HalfCheetah is a simulated benchmark, it captures key aspects of many practical robot control tasks, as they all rely on the same underlying design and physics principles.
This extension is motivated by prior work such as (KAN_RL), where the authors showed that KAN actor networks with significantly fewer trainable parameters achieved higher rewards compared to much larger MLPs when trained with the Proximal Policy Optimization (PPO) algorithm. PPO is one of the most widely used RL algorithms in practice, particularly in robotics, games, and continuous control tasks.
@@ 5.7.3. Hardware Performance
To evaluate the efficiency of KANs in practical deployment, we compare the hardware cost of the 8-bit quantized KAN actor against the 8-bit quantized MLP actor. The KAN actor was deployed using the KANELÉ framework, while the MLP actor was implemented with hls4ml (hls4ml_Fahim) using the Resource strategy for a fair baseline. We planned to have both models synthesized on a Xilinx xczu7ev-ffvc1156-2-e FPGA in out-of-context mode. However, the 8-bit MLP design exceeds the available FPGA resources, so its performance results are based on HLS estimates rather than the actual implementation. Table 7 summarizes the results. As can be seen, the 8-bit KAN actor achieves significantly lower resource utilization, latency, and power compared to the 8-bit MLP actor, underscoring the advantages of KANs for real-time reinforcement learning control tasks.

### v2

@@ 5.4. Comparison with Prior KAN-FPGA Literature
We benchmarked KANELÉ against the KAN FPGA implementation by Tran et al. (Tran et al., 2024) on the Moons, Wine, and Dry Bean datasets. The results, detailed in Table 4, demonstrate that KANELÉ offers a dramatic improvement in hardware efficiency and performance while maintaining or exceeding the accuracy of the previous work.
@@ 5.5. Comparison with hls4ml MLPerf Tiny
To understand the performance of KANELÉ on more complex datasets, we compare the framework to hls4ml (Borras et al., 2022) on the ToyADMOS dataset, part of the MLPerf Tiny benchmark suite. Other LUT-based neural networks have not attempted this benchmark, possibly due to its high complexity and/or non-classification-based training scheme. The results in Table 5 indicate that KANELÉ achieves substantially better performance than prior approaches in terms of both resource efficiency, latency, and power. These improvements highlight the potential of KANELÉ for future studies involving more complex datasets as well as tasks beyond classification. Specifically, KANELÉ eliminates the need for BRAM, LUTRAM, and DSPs, while reducing LUT usage by 41.7% and FF usage by 71.4% relative to hls4ml. In terms of performance, KANELÉ delivers $330\times$ higher throughput, $643\times$ lower latency and a $9,840\times$ reduction in energy per inference. These results establish KANELÉ as a highly efficient alternative to existing FPGA neural implementations, with strong potential for scaling to more complex datasets and tasks beyond classification.
@@ 5.7. Extension to Real-time Control Systems
To demonstrate that the KANELÉ paradigm extends well beyond the traditional supervised learning tasks typically studied in the LUT-based neural network community, we apply our framework to a reinforcement learning benchmark. Specifically, we evaluate on the HalfCheetah environment, a classic continuous control task in reinforcement learning (RL), most commonly accessed through the MuJoCo physics simulator via OpenAI Gym (now Gymnasium) (Todorov et al., 2012). The objective is to learn a policy that enables a simulated two-legged agent to run as fast and as stably as possible. This environment is widely used as a standard benchmark to compare RL algorithms and function approximators. While HalfCheetah is a simulated benchmark, it captures key aspects of many practical robot control tasks, as they all rely on the same underlying design and physics principles.
This extension is motivated by prior work such as (Kich et al., 2024), where the authors showed that KAN actor networks with significantly fewer trainable parameters achieved higher rewards compared to much larger MLPs when trained with the Proximal Policy Optimization (PPO) algorithm. PPO is one of the most widely used RL algorithms in practice, particularly in robotics, games, and continuous control tasks.
@@ 5.7.3. Hardware Performance
To evaluate the efficiency of KANs in practical deployment, we compare the hardware cost of the 8-bit quantized KAN actor against the 8-bit quantized MLP actor. The KAN actor was deployed using the KANELÉ framework, while the MLP actor was implemented with hls4ml (Fahim et al., 2021) using the Resource strategy for a fair baseline. We planned to have both models synthesized on a Xilinx xczu7ev-ffvc1156-2-e FPGA in out-of-context mode. However, the 8-bit MLP design exceeds the available FPGA resources, so its performance results are based on HLS estimates rather than the actual implementation. Table 7 summarizes the results. As can be seen, the 8-bit KAN actor achieves significantly lower resource utilization, latency, and power compared to the 8-bit MLP actor, underscoring the advantages of KANs for real-time reinforcement learning control tasks.